A radio disc jockey has 8 songs on this upcoming hour's playlist: 2 are rock songs, 2 are reggae songs, and 4 are country songs. The disc jockey randomly chooses the first song to play, and then she randomly chooses the second song from the remaining ones. What is the probability that both songs are rock songs? Write your answer as a fraction in simplest form.
Answer by jim_thompson5910(35256)


There are 2 rock songs out of 8 total. So the probability of picking a rock song is 2/8 = 1/4

After the first rock song is chosen, we have 2-1 = 1 rock song left out of 8-1 = 7 total. The probability of selecting another rock song is 1/7

Multiply the two probabilities: (1/4)*(1/7) = (1*1)/(4*7) = 1/28